### v2

# Yang-Mills extension of the Loop Quantum Gravity-corrected Maxwell equations

Abstract
In this paper, we endeavour to build up a non-Abelian formulation to describe the self-interactions of massless vector bosons in the context of Loop Quantum Gravity (LQG). To accomplish this task, we start off from the modified Maxwell equations with the inclusion of LQG corrections and its corresponding local $U(1)$ gauge invariance. LQG effects in the electromagnetic interactions have significant importance, as they might be adopted to describe the flight time of cosmic photons coming from very high-energy explosions in the Universe, such as events of Gamma-Ray Bursts (GRBs). These photons have energy-dependent speeds, indicating that the velocity of light in the vacuum is not constant. To carry out the extension from the Abelian to the non-Abelian scenario, we shall follow the so-called Noether current procedure, which consists in recurrently introducing self-interactions into an initially free action for vector bosons by coupling the latter to the conserved currents of a global symmetry present in the action of departure. In the end of the non-Abelianization process, the initial global symmetry naturally becomes local. Once the Yang-Mills system includes LQG correction terms, it becomes possible to analyze how quantum-gravity induced contributions show up in both the electroweak and the QCD sectors of the Standard Model, providing a set-up for phenomenological investigations that may bring about new elements to discuss Physics beyond the Standard-Model.
I Introduction
The consistent and precise description of quantum gravity is a big challenge and an important open problem in the theory of fundamental interactions. This quest is going on for more than 50 years, by adopting different scenarios. We point out that some of those try to achieve their goal by also seeking a grand-unified theory. In this framework, one of the most promising theories is Loop Quantum Gravity (LQG) Ashtekar:2004eh ; Ashtekar:2011ni ; Rovelli:2014ssa ; Rovelli:1997yv , whose main purpose is to understand fundamental physics at the Planck scale $l_{P}=10^{-35}$m, through the quantization of gravity and not via a grand-unification approach. The LQG formulation predicts that spacetime must be discrete, granular, at the Planck scale. More precisely, the granularity of spacetime is implemented by means of a mathematically structure referred to as spin foam Rovelli:2014ssa ; Rovelli:1997yv ; Thiemann:2001gmi ; Gambini:2011zz . Moreover, the fields present a background-independent description that manifests itself through an action that exhibits diffeomorphism invariance and is also non-perturbative. There are many ways to describe LQG; in the present article, we choose to work with the canonical method, which is described by Hamiltonians Thiemann:2001gmi . The formalism of LQG is based on the Ashtekar variables Ashtekar:1986yd , introduced in by 1986 Abhay Ashtekar in his formalism to describe General Relativity by an $SU(2)$ gauge connection.
Loop Quantum Gravity offers a consistent theoretical framework which extends its exploration to include phenomena that contribute to the understanding of what gravity is at a fundamental quantum scale. If we aim to detect possible traces of LQG in various events Amelino-Camelia:1997ieq ; Ellis:1999rz ; Amelino-Camelia:1999hpv , the most notable ones are: noise present in gravitational wave detectors Amelino-Camelia:2001dbf ; Amelino-Camelia:1999vks , neutral kaon systems Ellis:1995xd ; Huet:1994kr and the time-dependent energy of arrival of photons and neutrinos from distant sources Alfaro:2001rb ; Li:2022szn . The electromagnetic sector has an important contribution to this. It is expected that at the Planck scale (local) Lorentz invariance, a basic principle of General Relativity, be broken. However, achieving the energy scales necessary to test both Lorentz-invariance violation (LIV) and Planck-scale effects in terrestrial experiments is not feasible Amelino-Camelia:1997ieq ; Li:2022szn ; Gambini:1998it ; Amelino-Camelia:1996bln ; Melo:2024gxl . An alternative is to focus on the energy of Gamma-Ray Bursts (GRBs) and their enormous energy emitted, which may render possible to investigate the energy-dependent time of arrival of photons/neutrinos from distant sources. In this work, we are committed with the photon sector. The speed of light in a vacuum characterized by a granular space-time may assume the form
$v(E)\simeq c\left(1-\frac{E}{E^{\gamma}_{LIV}}\right),$ (1)
where $E^{\gamma}_{LIV}\approx 3.6\times 10^{17}$ GeV and is independent of the photon’s helicity. The high-energy scale of the $E^{\gamma}_{LIV}$-parameter supports the idea that GRBs are the most suitable candidates to investigate this parameter  Shao:2009bv ; Zhang:2014wpb ; Xu:2016zxi ; Xu:2016zsa ; Amelino-Camelia:2016ohi ; Amelino-Camelia:2017zva ; Xu:2018ien ; Liu:2018qrg ; Li:2020uef ; Zhu:2021pml ; Chen:2019avc .
Yang-Mills (Y-M) theories are widely adopted by virtue of their effectiveness in describing the physics of fundamental interactions with the remarkable attribute of naturally opening up the path to unification. In LQG, it could not be different; there are already several works that present such a formulation as a result of the complete theory such as Thiemann:2001gmi ; Rovelli:1997yv ; Gambini:1996ik . In this paper, our goal is not to build up a non-Abelian theory in the full-fledged LQG. It is important to stress that our actual purpose is to start off from the already-known set of Maxwell’s equations including LQG effects and, through the so-called the Noether’s procedure to write down self-interactions among spin-1 fields, we get to the non-Abelian extension of the LQG Maxwell’s equations, which we adopt the LQG-corrected Yang-Mills field equations. We intend thereby to describe a way how quantum gravitational effects, as dictated by LQG, may be brought to the Yang-Mills theory. This helps us in pursuing an investigation of the physics of self-interacting massless vector bosons in presence of tiny quantum gravity effects with the LQG signature. Effects of quantum gravity on both photons and neutrinos may be studied, as it can be seen in Li:2022szn . Inspired by this reference paper, our work, by putting together LQG and Yang-Mills, may offer a framework to study how the electroweak and QCD sectors get quantum-gravitational corrections, contributing to an interesting phenomenological analysis and raising the possibility of new hidden physics. The working method adopted in this paper is, as already mentioned, the so-called Noether method Deser:1969wk ; VanNieuwenhuizen:1981ae . It is not a very popular method, but is very effective in describing this type of problem and is widely found in the literature of supersymmetry and supergravity.
In this paper, we are adopting the natural system of units, $c=\hbar=1$ and the signature of Minkowski metric is (+,,,)(+,-,-,-)( + , - , - , - ). Our work is outlined as follows. In Sec. II, we briefly introduce the electromagnetic sector in LQG; for this, we consider the Hamiltonian of the theory and the four modified Maxwell equations in the context of Loop Quantum Gravity. In Section  III, we show the details how, from the Abelian electromagnetic sector with its Hamiltonian and equations of motion, we can obtain a non-Abelian theory of self-interactions for massless vector (gauge) bosons. Finally, in Sec. IV, we present our final comments and discussions.
II LQG Effects in the Electromagnetic Sector
The Hamiltonian formulation of the effects of LQG in the electromagnetic theory was carried out in the work of Ref.  Alfaro:2001rb . This construction was done by considering the spacetime as a manifold $M$ whose topology is a diffeomorphism for $\Sigma\times\mathbb{R}$, where $\Sigma$ is a Riemmannian-3 manifold. The LQG formalism is based on the Ashtekar variables described for a non-compact gauge group, $SL(2,\mathbb{C})$, where the indices $a,b,c$ stand for the space coordinates and $i,j,k$ are their indices assigned to the generators of $su(2)$. The gauge connection is represented as $(A_{a}^{i}=\Gamma_{a}^{i}-iK_{a}^{i})$, with $\Gamma_{a}^{i}$, torsion-free, is a function of the co-triads $e^{i}_{a}$, which in turn are defined as a function of the three-metric $q_{ab}=e^{i}_{a}e^{i}_{b}$. The term $K_{a}^{i}$ is the extrinsic curvature of the Riemmannian 3-manifold, $\Sigma$, and is related to the co-triads according to the relation $K_{ab}=sgn[\mbox{det}(e_{c}^{j})]K_{a}^{i}e_{b}^{i}$. The other variable part of Ashtekar’s formalism is the canonical pair of the conjugate moment, namely, the electric field of the form $E_{i}^{a}=\frac{1}{2}\epsilon^{abc}\epsilon_{ijk}e_{b}^{j}e_{c}^{k}$. The Hamiltonian of the electromagnetic sector modified by LQG is built up based on these considerations, and it takes the form:
$\begin{split}H_{LQG}=\frac{1}{Q^{2}}\int d^{3}x\Biggl{\{}\left[1+\theta_{7}\left(\frac{l_{P}}{\mathcal{L}}\right)^{2+2\Upsilon}\right]\frac{1}{2}(\underline{\vec{B}}^{2}+\underline{\vec{E}}^{2})\\ +\theta_{3}l_{P}^{2}(\underline{B}^{a}\nabla^{2}\underline{B}_{a}+\underline{E}^{a}\nabla^{2}\underline{E}_{a})+\theta_{2}l_{P}^{2}\underline{E}^{a}\partial_{a}\partial_{b}\underline{E}^{b}\\ +\theta_{8}l_{P}[\underline{\vec{B}}\cdot(\nabla\times\underline{\vec{B}})+\underline{\vec{E}}\cdot(\nabla\times\underline{\vec{E}})]\\ +\theta_{4}\mathcal{L}^{2}l_{P}^{2}\left(\frac{\mathcal{L}}{l_{p}}\right)^{2\Upsilon}(\underline{\vec{B}}^{2})^{2}+...\Biggl{\}},\end{split}$ (2)
where $Q^{2}$ is the coupling constant of electromagnetism, $l_{p}\approx 1.6\times 10^{-35}m$ is the Planck length. The characteristic length, $\mathcal{L}$, satisfies the condition $l_{P}\ll\mathcal{L}\leq\lambda$. The parameter $\lambda$ is the de Broglie wavelength, and the characteristic length, $\mathcal{L}$, has a maximum value at momentum $k$ whenever $\mathcal{L}=k^{-1}$. Another parameter appearing in the equation is $\Upsilon$, which is given by the order of the contribution of the gravitational connection to the expected value and it may be determined by the phenomenological analysis of a selected event. That is, it may depend on the helicity of the particle under consideration Ellis:1999uh ; Ellis:1999yd ; Ellis:1999sd . The $\theta_{i}$’s are non-dimensional parameters of order one or are extremely close to zero Alfaro:2001rb ; Alfaro:2002ya ; finally, $a,b$ are spatial tensor indices. From now on, we replace the space indices by adopting a vector notation, and we eliminate the underline that refers to the canonical pairs for the electromagnetic sector in all quantities. From the Eq. (2), we get the field equations as follows below:
$\nabla\cdot\vec{E}=0,$ (3)
$\begin{split}A_{\gamma}(\nabla\times\vec{B})-\frac{\partial\vec{E}}{\partial t}+2l_{P}^{2}\theta_{3}\nabla^{2}(\nabla\times\vec{B})-2\theta_{8}l_{P}\nabla^{2}\vec{B}\\ +4\theta_{4}\mathcal{L}^{2}\left(\frac{\mathcal{L}}{l_{P}}\right)^{2\Upsilon_{\gamma}}l_{P}^{2}\nabla\times(\vec{B}^{2}\vec{B})=0,\end{split}$ (4)
$\nabla\cdot\vec{B}=0,$ (5)
$\begin{split}A_{\gamma}(\nabla\times\vec{E})+\frac{\partial\vec{B}}{\partial t}+2l_{P}^{2}\theta_{3}\nabla^{2}(\nabla\times\vec{E})-\\ 2\theta_{8}l_{P}\nabla^{2}\vec{E}=0,\end{split}$ (6)
with
$A_{\gamma}=1+\theta_{7}\left(\frac{\ell_{p}}{\mathcal{L}}\right)^{2+2\Upsilon}.$ (7)
The Eq. (2) is an infinite expression. In the literature, it is common to truncate it at the first order in the non-linear term in the magnetic field. When the field equations are calculated, this non-linear term contributes to Eq. (4). The main quantities of the electromagnetic theory have already been calculated and analyzed previously, as it can be seen in Alfaro:2001rb ; Li:2020uef . In the present work, such quantities are not of primary interest; the modified Maxwell equations in the context of LQG as well as its corresponding Hamiltonian are the key equations for the development of the present contribution.
III The Noether Procedure as our path towards the non-Abelian scenario
To formulate a non-Abelian theory for massless spin-1 particles, one may follow different approaches. In this paper, we pick out the so-called Noether method Deser:1969wk ; VanNieuwenhuizen:1981ae , which essentially consists in generating self-interactions among vector fields through the variation of the original free Lagrangian and proposing self-interactions by coupling these vector fields with the Noether current of their initially global. To realize this program, we start off from an N-plet of vector fields (at the end of the procedure, they will become our Yang-Mills gauge bosons) in an arbitrary representation of the symmetry group as given below:
$A^{m},m=1,2,...,N\ \ \ \ \ \ \ \phi^{n},n=1,2,...,N$ (8)
in N-dimensions, we can represent these fields in a Lie group that transforms as
$\phi_{m}^{\prime}=R_{mn}\phi_{n},$ (9)
and
$A_{m}^{\prime}=R_{mn}A_{n}.$ (10)
Consider an auxiliary function of the Lie group, which can be expressed in infinitesimal form as Peskin:1995ev :
$R_{mn}=(e^{iw_{h}G_{h}})_{mn}\sim\delta_{mn}+iw_{h}(G_{h})_{mn}+O(w^{2}),$ (11)
similarly, the fields can also be expressed in a form
$\delta A_{m}=iw_{h}(G_{h})_{mn}A_{n},$ (12)
and
$\delta\phi_{m}=iw_{h}(G_{h})_{mn}\phi_{n},$ (13)
note that $(G_{h})_{mn}$ are the generators in the adjoint representation; they are written in terms of the structure constants as given below:
$[G_{h},G_{i}]=if_{hij}G_{j}.$ (14)
We are working in the general case of an $SU(N)$ gauge group. So, $w_{h}$ stands for the parameters of the $SU(N)$ transformations, whereas the 3-index symbol $f_{hij}$ represents the structure constants of $SU(N)$. To initiate the Noether procedure, we start off with the free Lagrangian of the theory. To get this Lagrangian, we perform a Legendre transformation by taking Eq. (2) and writing
$\begin{split}L_{LQG}=\frac{1}{Q^{2}}\int d^{3}x\Biggl{\{}\left[1+\theta_{7}\left(\frac{l_{P}}{\mathcal{L}}\right)^{2+2\Upsilon}\right]\frac{1}{2}(\vec{E}^{2}-\vec{B}^{2})\\ +\theta_{3}l_{P}^{2}(\vec{B}\nabla^{2}\vec{B}+\vec{E}\nabla^{2}\vec{E})+\theta_{2}l_{P}^{2}\vec{E}\cdot[\nabla\cdot(\nabla\vec{E})]\\ +\theta_{8}l_{P}[\vec{B}\cdot(\nabla\times\vec{B})+\vec{E}\cdot(\nabla\times\vec{E})]\\ +\theta_{4}\mathcal{L}^{2}l_{P}^{2}\left(\frac{\mathcal{L}}{l_{p}}\right)^{2\Upsilon}(\vec{B}^{2})^{2}\Biggl{\}},\end{split}$ (15)
Using the variational principle, we have
$\delta S=\int d^{4}x\ \delta L_{LQG}=0.$ (16)
By considering the modified Maxwell equations in the context of LQG, it is possible to extract the (on-shell) conserved currents. Notice that these currents carry the index ($h$) of the adjoint representation of the symmetry group, while the fields carry the index of an arbitrary representation, ($m$). Therefore, invariant self-interaction can only occur if these indices coincide, i.e., $m=h$. To allow for self-interaction between fields and currents, it is necessary that the fields, which were previously in an arbitrary representation, be placed in the adjoint representation, as the currents are. The matrix elements of the generators in the adjoint representation are written in terms of the structure constants of the symmetry group:
$(G_{h})_{ij}=-if_{hij},$ (17)
so the spatial current is
$\begin{split}\vec{j}_{h}=-f_{hij}\{A_{\gamma}[(\vec{E}_{i}\cdot\phi_{j})-(\vec{B}_{i}\times\vec{A}_{j})]-\theta_{3}l_{P}^{2}[(\vec{B}_{i}\times\\ \nabla^{2}\vec{A}_{j})-(\nabla^{2}\vec{B}_{i}\times\vec{A}_{j})-(\nabla^{2}\vec{E}_{i}\cdot\phi_{j})-(\vec{E}_{i}\cdot\nabla^{2}\phi_{j})]\\ -\theta_{8}l_{P}[2(\nabla^{2}\vec{A}_{i}\times\vec{A}_{j})-(\vec{E}_{i}\times\vec{\dot{A}}_{j})-(\nabla\times\vec{E}_{i})\phi_{j}]\\ +4\theta_{4}\mathcal{L}^{2}l_{P}^{2}\left(\frac{\mathcal{L}}{l_{p}}\right)^{2\Upsilon}(B^{2}\cdot\vec{B}_{i}\times\vec{A}_{j})\},\end{split}$ (18)
and the temporal current as
$\begin{split}j_{h}^{0}=-f_{hij}\{A_{\gamma}(\vec{E}_{i}\cdot\vec{A}_{j})+\theta_{3}l_{P}^{2}[(\nabla^{2}\vec{E}_{i}\cdot\vec{A}_{j})+\\ (\vec{E}_{i}\cdot\nabla^{2}\vec{A}_{j})]-2\theta_{8}l_{P}(\nabla\times\vec{E}_{i})\cdot\vec{A}_{j}\}.\end{split}$ (19)
By coupling the Noether currents to the fields , we go over from the free regime to one that exhibits self-interaction among the massless vector fields. We then introduce the coupling between the currents and the fields to get a new Lagrangian, no more free. This yields:
$L_{LQG}^{1}=L_{LQG}-l\phi_{h}j_{h}^{0}-l\vec{A_{h}}\vec{j}_{h},$ (20)
the parameter $l$ is the coupling constant for the self-coupling introduced above. The development is not yet complete, because it is necessary to remove any dependence of the partial Lagrangian on space-time derivatives of the vector fields. This means that we have repeat the same procedure again. Due to the currents presenting long expressions, we chose not to present them here. Let us consider the following infinitesimal transformations for the fields
$\delta A_{h}=iw_{k}(G_{k})_{hl}A_{l}=w_{k}f_{khl}A_{l},$ (21)
and
$\delta\phi_{h}=iw_{k}(G_{k})_{hl}\phi_{l}=w_{k}f_{khl}\phi_{l}.$ (22)
Again, we apply the principle of least action, but now the Lagrangian used is the one in Eq. (20)
$\delta S=\int d^{4}x\ \delta L_{LQG}^{1}=0.$ (23)
The process will be the same, but with an additional step. Before identifying the new conserved currents, we must consult the new field equations derived from the Lagrangian of the previous step. This is so because the new identified currents must be on-shell. This is an essential point. From the Lagrangian, we can derive new Gauss’ electricity and Ampère-Maxwell equations. This allows us to find the currents to be coupled to the fields to once again insert in the Lagrangian. Each current is coupled to a field and depends on a coupling parameter $l^{\prime}$, which represents the coupling constant for this new term
$L_{LQG}^{2}=L_{LQG}^{1}-l^{\prime}\phi_{k}j_{k}^{0(1)}-l^{\prime}\vec{A_{k}}\vec{j}_{k}^{(1)},$ (24)
The Eq. (24) is the last Lagrangian expression obtained, and it is no longer necessary to re-calculate other currents. The reason being that the expression of the new Lagrangian no longer depends on derivatives of the fields. Should the same steps be repeated, we would obtain the same expressions for the currents as before. From Eq. (24), we can obtain the field equations, namely, Gauss’ electricity and Ampère-Maxwell. We set $l=\frac{1}{4}g$ and $l^{\prime}=\frac{1}{2}g$ and from these two equations of motion, it is possible to define the equations for the electric and magnetic field equations in terms of the potentials in the Yang-Mills formalism with LQG contributions:
$\begin{split}\vec{E}_{h}=A_{\gamma}[-\nabla\cdot\phi_{h}-\vec{\dot{A}}_{h}+gf_{hij}\phi_{i}\vec{A}_{j}]-\frac{3g}{2}\theta_{3}l_{P}^{2}f_{hij}\cdot\\ [\vec{A}_{i}(\nabla^{2}\phi_{j})+(\nabla^{2}\vec{A}_{i})\cdot\phi_{j}]+\theta_{3}l_{P}^{2}f_{hij}[(\nabla\vec{A}_{i})\cdot(\nabla\phi_{j})]\\ +\frac{g}{2}\theta_{8}l_{P}f_{hij}(\vec{A}_{i}\times\vec{\dot{A}}_{j}),\end{split}$ (25)
and
$\begin{split}\vec{B}_{h}=A_{\gamma}[(\nabla\times\vec{A}_{h})+\frac{g}{2}f_{hij}(\vec{A}_{i}\times\vec{A}_{j})]+2l_{P}^{2}\theta_{3}\nabla^{2}\vec{B}_{h}\\ -\frac{g}{2}\theta_{3}l_{P}^{2}f_{hij}\cdot[\nabla^{2}(\vec{A}_{i}\times\vec{A}_{j})-[\vec{A}_{i}\times(\nabla^{2}\vec{A}_{j})]]+\\ \frac{g}{2}\theta_{8}l_{P}f_{hij}[(\vec{\dot{A}}_{i})\cdot\phi_{j}]+2\theta_{4}\mathcal{L}^{2}\left(\frac{\mathcal{L}}{l_{P}}\right)^{2\Upsilon_{\gamma}}l_{P}^{2}[2(\vec{B}^{2}\vec{B}_{h})\\ +3gf_{hij}(\vec{A}_{i}\times\vec{A}_{j})\cdot B^{2}].\end{split}$ (26)
If we were working with the electromagnetism of the Standard Model of particle physics, all the procedures described so far would be sufficient to obtain all the non-Abelian Maxwell equations. Simply taking the curl of Eq. (25) and the divergence of Eq. (26) would be sufficient. In electrodynamics with LQG effects, this is not possible, because the Faraday-Lenz equation is modified and, by doing this and taking the Abelian limit, we do not obtain again the Eqs. (5), (6) that should be obtained. As it can be consulted in the literature Sorokin:2021tge ; Gaete:2014nda , the non-Maxwellian extensions of electrodynamics does not modify the Faraday-Lenz equation, since together with Gauss’law for magnetism, these equations are obtained from the Bianchi Identity. The electrodynamics presented in this paper modifies Eq. (6), which allows us to conclude that the Bianchi Identities are modified. It is therefore necessary to follow another path to attain these two equations; one possible way consists in getting them from the Hamilton-Jacobi equations. By performing a Legendre transformation on the Lagrangian Eq. (24) to find its corresponding Hamiltonian, we can derive the equations for the electric field, $E$, and the vector potential, $A$:
$\frac{\partial H_{LQG}^{2}}{\partial\vec{E}_{h}}=-\frac{\partial\vec{A}_{h}}{\partial t},$ (27)
and
$\frac{\partial H_{LQG}^{2}}{\partial\vec{A}_{h}}+\frac{\partial H_{LQG}^{2}}{\partial(\nabla\times\vec{A}_{h})}=\frac{\partial\vec{E}_{h}}{\partial t}.$ (28)
This yields four Maxwell-type equations in the non-Abelian version with LQG contributions :
$2A_{\gamma}[\nabla\cdot\vec{E}_{h}+gf_{hij}\vec{E}_{i}\cdot\vec{A}_{j}]+g\theta_{3}l_{P}^{2}f_{hij}[2(\nabla^{2}\vec{E}_{i})\cdot\vec{A}_{j}+\vec{E}_{i}\cdot(\nabla^{2}\vec{A}_{j})+\nabla^{2}(\vec{A}_{i}\cdot\vec{E}_{j})-$
 $\;2\nabla[(\nabla\vec{A}_{i})\cdot(\nabla\phi_{j})]]+g\theta_{8}l_{P}f_{hij}[\nabla\cdot(\vec{A}_{i}\times\vec{\dot{A}}_{j})+\vec{A}_{i}\cdot(\nabla\times\vec{E}_{j})]=0,$ (29)
$A_{\gamma}[\nabla\times\vec{B}_{h}-gf_{hij}\phi_{i}(\vec{\dot{A}}_{j})-gf_{hij}\phi_{i}\vec{E}_{j}+gf_{hij}(\vec{B}_{i}\times\vec{A}_{j})]+\theta_{3}l_{P}^{2}\{2\nabla^{2}(\nabla\times\vec{B}_{h})-gf_{hij}[(\nabla^{2}\vec{E}_{i})\phi_{j}+$
$\;\frac{1}{2}\nabla^{2}(\phi_{i}\vec{E}_{j})+[(\nabla^{2}\vec{B}_{i})\times\vec{A}_{j}]-\frac{1}{2}\phi_{i}\nabla^{2}(\vec{\dot{A}}_{j})+\frac{1}{2}\vec{E}_{i}(\nabla^{2}\phi_{j})-\frac{1}{2}(\vec{\dot{A}}_{i})\cdot(\nabla^{2}\phi_{j})-\frac{1}{2}\nabla^{2}(\vec{B}_{i}\times\vec{A}_{j})-$
$\;\frac{1}{2}[\vec{B}_{i}\times(\nabla^{2}\vec{A}_{j})]-\nabla^{2}[\phi_{i}(\vec{\dot{A}}_{j})]]\}-\theta_{8}l_{P}\{2\nabla^{2}\vec{B}_{h}-\frac{g}{2}f_{hij}[\partial_{t}(\vec{A}_{i}\times\vec{A}_{j})-[(\vec{\dot{E}}_{i})\times\vec{A}_{j}]-2\nabla^{2}(\vec{A}_{i}\times\vec{A}_{j})-$
$\;\nabla[\nabla\cdot(\vec{A}_{i}\times\vec{A}_{j})]+(\nabla\times\vec{E}_{i})\phi_{j}]\}+2\theta_{4}\mathcal{L}^{2}\left(\frac{\mathcal{L}}{l_{P}}\right)^{2\Upsilon_{\gamma}}l_{P}^{2}\nabla\times\left[2(\vec{B}^{2}\vec{B}_{h})+3gf_{hij}(\vec{A}_{i}\times\vec{A}_{j})\cdot B^{2}\right]=\frac{\partial\vec{E}_{h}}{\partial t},$ (30)
$A_{\gamma}[(\nabla\cdot\vec{B}_{h})+gf_{hij}\vec{A}_{i}\cdot\vec{B}_{j}]-\frac{g}{2}\theta_{3}l_{P}^{2}f_{hij}\nabla\cdot[(\vec{A}_{i}\times(\nabla^{2}\vec{A}_{j})-3\nabla^{2}(\vec{B}_{i}\times\vec{A}_{j})]+\frac{g}{2}\theta_{8}l_{P}f_{hij}\nabla\cdot[(\vec{\dot{A}}_{i})\cdot\phi_{j}]+$
 $\;9g\theta_{4}\mathcal{L}^{2}\left(\frac{\mathcal{L}}{l_{P}}\right)^{2\Upsilon_{\gamma}}l_{P}^{2}f_{hij}\nabla\cdot[(\vec{A}_{i}\times\vec{A}_{j})\cdot B^{2}]=0,$ (31)
$A_{\gamma}[(\nabla\times\vec{E}_{h})-gf_{hij}\phi_{i}\vec{B}_{j}+gf_{hij}(\vec{A}_{i}\times\vec{E}_{j})]+\theta_{3}l_{P}^{2}\nabla^{2}(\nabla\times\vec{E}_{h})-\frac{g}{2}\theta_{3}l_{P}^{2}f_{hij}[\vec{B}_{i}\cdot(\nabla^{2}\phi_{j})-$
$\;(\nabla^{2}\vec{E}_{i}\times\vec{A}_{j})]-\theta_{8}l_{P}[\nabla^{2}\vec{E}_{h}+gf_{hij}\nabla\cdot(\vec{E}_{i}\cdot\vec{A}_{j})-\frac{g}{2}f_{hij}\nabla\times(\vec{A}_{i}\times(\vec{\dot{A}}_{j}))]=-\frac{\partial\vec{B}_{h}}{\partial t}.$ (32)
When considering the four non-Abelian Maxwell-type equations for massless spin-1 vector bosons in the context of LQG effects, it is important to notice that when the generators of the group $SU(N)$ become trivial ($G_{h}=0$), the structure constants of the group are taken $f_{hij}=0$, we re-obtains the LQG-corrected Maxwell equations, (3), (4), (5), and (6). Each term in $\theta^{i}$ contributes independently to the structure constants, and the term $\theta_{4}$, coupled with the nonlinear magnetic field, modifies the new magnetic field equations. Previously, the Eq. (31), known as Gauss Magnetism, did not include magnetic monopoles. However, this is no longer true in the non-Abelian formulation, as terms $\theta_{3}$, $\theta_{8}$, and $\theta_{4}$ independently contribute to this equation. It is conceivable, based on this interpretation, that magnetic monopoles might exist as the Planck scale, possibly associated to these parameters.
IV Conclusive Considerations
In this work, all our efforts consist in formulating a non-Abelian action for massless spin-1 fields in the context of Loop Quantum Gravity. We know that Yang-Mills gauge theories give a very complete description of fundamental interactions: this is the case of QCD and the unified electroweak model based on the $SU(2)\times U(1)$-symmetry.
To formulate the theory, we first transformed the free Hamiltonian into a free Lagrangian by means of a Legendre transformation. By applying the variational principle, we were able to obtain the time and space components of the Noether currents. Next, we introduced these currents coupled to the vector fields into the free Lagrangian to generate self-interactions. The procedure was recursively repeated until we obtained another self-interaction current, which, when coupled to the vector fields and inserted into the step-by-step corrected Lagrangian stopped the process. Additionally, by deriving new motion equations from the new Lagrangian, we obtained the Gauss Electricity and Ampère-Maxwell equations. Furthermore, we were able to define the Electric and Magnetic fields in $SU(N)$. However, unlike the Abelian Maxwell equations, we could not obtain the last two equations, namely, Gauss magnetic and Faraday-Lenz equations. To resolve this, we obtained a Hamiltonian from the final Lagrangian once again, and from the Hamilton-Jacobi motion equations, we derived the last two equations.
The four modified Maxwell equations in the context of LQG and non-Abelian present significant results, as the dimensionless parameters of the LQG contribute to all equations. In the Abelian case, this does not happen, it is still possible to notice that the nonlinear parameter of the magnetic field in the Hamiltonian/Lagrangian is present in both equations, in which the magnetic field is the core of the discussion. Furthermore, this non-linear term, together with the LQG parameters and the other parameters that we have already mentioned, is present in the magnetic monopole equation. So, it is possible that on the Planck scale, these parameters can together yield magnetic monopole solutions. It is important to remark that we have not worked in the covariant formalism. The idea is to express in a manifest way field strengths and the gauge potentials interact. The developments proposed in this paper may open up the way for future developments of LQG effects in unexplored areas of physics, such as the investigation of LQG effects in both the electroweak theory and QCD. It is important to notice that effects arising from quantum gravitation may be found in these theories at high-energy scales. In particular, considering the electroweak sector, in our next project, we propose to pursue an investigation on the anomalous gauge-boson couplings - more specifically the anomalous (neutral) photon/Z-boson 3- and 4-vertices - searched for in the ATLAS and CMS experiments of the LHC, having in program to get new constraints on the LQG parameters from accelerator physics.
V Acknowledgements
G.L.L.W.L. acknowledges financial support of the
Fundação Carlos Chagas Filho de Amparo à Pesquisa
do Estado do Rio de Janeiro (FAPERJ), Grant No. E-
26/202.437/2024. We are also grateful to P. A. Lima
Mourão for valuable discussions on the LQG-corrected
Maxwell equations in presence of a background magnetic field.